how many liters of pure sale must be added to 15 liters of a 40% salt solution to obtain a 60% salt solution?
Answer by ankor@dixie-net.com(22740)

how many liters of pure salt must be added to 15 liters of a 40% salt solution to obtain a 60% salt solution?
:
let s = amt of pure salt required
the resulting amt will be (s+15)
:
.4(15) + s = .60(s + 15)
6 + s = .6s + 9
s - .6s = 9 - 6
.4s = 3
s = 3/.4
s = 7.5 liters of salt